# Investigating the Event-Shape Methods in Search for the Chiral Magnetic Effect in Relativistic Heavy Ion Collisions

### v4

@@ Abstract
The Chiral Magnetic Effect (CME) is a phenomenon in which electric charge is separated by a strong magnetic field from local domains of chirality imbalance and parity violation in quantum chromodynamics (QCD). The CME-sensitive observable, charge-dependent three-point azimuthal correlator $\Delta\gamma$, is contaminated by a major physics background proportional to the particle’s elliptic flow anisotropy $v_{2}$. Event-shape engineering (ESE) binning events in dynamical fluctuations of $v_{2}$ and event-shape selection (ESS) binning events in statistical fluctuations of $v_{2}$ are two methods to search for the CME by projecting $\Delta\gamma$ to the measured anisotropy $v_{2}=0$ intercept. We conduct a systematic study of these two methods using physics models as well as toy model simulations. It is observed that the ESE method fulfills the general premise of measuring the CME but is statistically hungry. It is found that the intercept from the ESS method depends on the details of the event content, such as the mixtures of background-contributing sources, because of statistical fluctuations of intertwining variables used in the method, and is thus not practically useful to measure the CME.
@@ I Introduction
The focus of our study is placed on the effects of the event classifying variables on the intercept under various circumstances of the simulated events. We note, however, that the $v_{2}$ measured in experiments contains not only the flow anisotropy but also the so-called nonflow correlations, dependent of analysis methods. Nonflow correlations refer to those unrelated to the collision geometry common to the entire event, such as resonance decays and (di-)jet correlations [37, 38, 39, 40]. While we touch upon the nonflow issue in this article, a full and thorough discussion of nonflow effects on the ESE and ESS methods is outside the scope of the present study.
@@
II The Event-Shape Techniques
@@
Since the ESE method selects events according to dynamical fluctuations of the $\left\langle v_{2}\right\rangle$ of the POIs, decoupled from the $q_{2}^{2}$ computation, it is guaranteed that the average resonance $\left\langle v_{\rm 2,{\rm res}}\right\rangle$, or generally the elliptic anisotropies of all elements in the event including all CME background sources, are proportional to $\left\langle v_{2}\right\rangle$. Note that the final-state $\left\langle v_{2}\right\rangle$ contains contributions not only from primordial particles, but also from all decay products of resonances and clusters. Since the averages $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ and $\left\langle v_{2}\right\rangle$ in those $q_{2}$-selected event classes are all connected to the initial geometry, they are all proportional to each other. This is illustrated in the cartoon in Fig. 1 (left), where the three ellipses illustrate event selections by $q_{2}$ and the hallow circles are the average $\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ in each $q_{2}$-selected event class. In the ESE analyses for CME searches [29, 30], one analyzes $\Delta\gamma$ as a function of $\left\langle v_{2}\right\rangle$ in each event class and projects the $\Delta\gamma$ measurement to zero $\left\langle v_{2}\right\rangle$. The ESE method guarantees that the elliptic flows of the background sources also vanish at zero $\left\langle v_{2}\right\rangle$. In other words, a linear projection to $\left\langle v_{2}\right\rangle=0$ would be equivalent to a projection to $\left\langle v_{\rm 2,{\rm res}}\right\rangle=0$, where the CME backgrounds are zero. Since all those events are from a narrow centrality bin where the CME signal does not vary much, the intercept of the linear projection is presumably the CME signal. Note that there may still be complications from nonflow effects [47] depending on how $\Delta\gamma$ and $\left\langle v_{2}\right\rangle$ are measured, which is outside the scope of the present work.
@@
Since the CME background comes from the elliptic anisotropies imprinted in the POIs, it is interesting to select events with an event-shape ellipticity $v_{2}^{\rm obs}\equiv\left\langle\cos 2(\phi-\psi)\right\rangle$ calculated by POIs themselves (where the EP $\psi$ is determined in another kinematic region away from the POIs) and examine the $\Delta\gamma$ observable as a function of $v_{2}^{\rm obs}$. This method selects on the event-by-event elliptic-shape quantity (particle emissin pattern) $v_{2}^{\rm obs}$ which can fluctuate even to negative values because of large statistical fluctuations, the intercept at $v_{2}^{\rm obs}=0$, more sensitive to the CME, can be well determined from data. This event-shape (ES) analysis was performed by STAR [24] and the intercept of the linear fit is consistent with zero with the then-available statistics. However, the backgrounds are not determined by $v_{2}^{\rm obs}$ but by $\left\langle v_{\rm 2,{\rm res}}\right\rangle$. Statistically fluctuated $v_{2}^{\rm obs}=0$ does not necessarily guarantee $\left\langle v_{\rm 2,{\rm res}}\right\rangle=0$. This is illustrated by the cartoon in Fig. 1 (center). Although $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ is proportional to $\left\langle v_{2}\right\rangle$ with an unbiased event selection, the strongly (maximally) biased selection of events by $v_{2}^{\rm obs}$ will obscure the relationship between the averages $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ and the $v_{2}^{\rm obs}$ in each bin of $v_{2}^{\rm obs}$ of those events, as indicated by the hollow box in Fig. 1 (center). Because of the overall positive means $\left\langle v_{2}^{\rm obs}\right\rangle=\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ over the entire event sample, the linear projection will give a positive intercept of $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ at $v_{2}^{\rm obs}=0$. In other words, with finite positive average $\left\langle v_{2}\right\rangle$ in heavy ion collisions, $\left\langle v_{\rm 2,{\rm res}}\right\rangle>0$ at $v_{2}^{\rm obs}=0$ with events binned in $v_{2}^{\rm obs}$. This has been verified by model study [48]. Thus, positive residual background remains in the linear fit intercept in the analysis of $\Delta\gamma$ as a function of $v_{2}^{\rm obs}$ [24], although the background is significantly reduced compared to that in the inclusive $\Delta\gamma$ measurement. Ideally, in the ES analysis, one would want to select events with the $v_{2}^{\rm obs}$ values of all resonances (background sources) to be zero, however, such selection is experimentally insurmountable.
@@
It has been recently proposed by Xu et al. [36] to select events according to the $\hat{q}_{2,\textsc{poi}}^{2}$ of the POIs, and study the $\Delta\gamma$ as a function of $v_{2}$ of the POIs. This is similar in spirit to the ES method above where the events are selected by $v_{2}^{\rm obs}$ and the $\Delta\gamma$ is examined as a function of the same $v_{2}^{\rm obs}$ variable. Here, the selection variable $\hat{q}_{2,\textsc{poi}}^{2}$ and the $v_{2}$ variable are different, but related as they are computed by the same POIs. The $\hat{q}_{2}^{2}$ variable is an event-by-event quantity connected to two-particle cumulant anisotropy (cf Eq. 5), whereas $v_{2}$ can be measured with respect to $\psi$ from another kinematic region like in the ES method. (Of course $v_{2}$ can also be computed by the two-particle cumulant method, in which case it is more directly connected to $\hat{q}_{2}^{2}$.) The variations in the event-by-event $\hat{q}_{2,\textsc{poi}}^{2}$ quantity, within a given narrow centrality bin, are mainly from statistical fluctuations, just like the $\hat{q}_{2}^{2}$ in the ESE method. However, unlike ESE, the $v_{2}$ variable in those $\hat{q}_{2,\textsc{poi}}^{2}$-selected events by ESS are related to $\hat{q}_{2,\textsc{poi}}^{2}$ itself and are calculated from the same POIs, so the variations in $v_{2}$ is also primarily of statistical nature. (Note, we simply use $v_{2}$ here, instead of the $\left\langle v_{2}\right\rangle$ for ESE, to indicate that the $v_{2}$ in ESS is of statistical nature although it is an “average” $v_{2}$ in events selected by $\hat{q}_{2,\textsc{poi}}^{2}$.) This technique selects on the statistical fluctuations in event shape, not on the dynamical fluctuations in $\left\langle v_{2}\right\rangle$ like in the ESE method, and it is referred to as the Event-Shape-Selection (ESS) method [36]. The most important difference in the ESS method from the ESE method is that the $\hat{q}_{2}^{2}$ is computed from the POIs in the former but in the later the momentum space where to compute $\hat{q}_{2}^{2}$ is displaced from that of the POIs’. In both approaches, one groups events (within a given narrow centrality bin) according to $\hat{q}_{2}^{2}$ and studies the $\Delta\gamma$ variable as a function of the $v_{2}$ of the POIs. It is worthy to emphasize again that, while the event selection variable in both methods are dominated by statistical fluctuations event-by-event, these statistical fluctuations are canceled in $v_{2}$ in the ESE method because the calculations of $\hat{q}_{2}^{2}$ and $v_{2}$ use different sets of particles, but do not cancel (rather strongly correlated) in the $v_{2}$ in the ESS method because the calculations of $\hat{q}_{2,{\rm pair}}^{2}$ and $v_{2}$ use the same POIs.
@@
One can study the $\Delta\gamma$ as a function of $v_{2}$ of single POI particles (or $v_{2,{\rm pair}}$ of pairs of POI particles) selecting events on $\hat{q}_{2,\textsc{poi}}^{2}$ of single POIs (or $\hat{q}_{2,{\rm pair}}^{2}$ of pairs of POIs). It is found by the Anomalous-Viscous Fluid Dynamics (avfd) model study that, out of the four combinations, it is the best to use the combination of pair $\hat{q}_{2,{\rm pair}}^{2}$ and single $v_{2,{\rm single}}$ [36], because the linear intercept of $\Delta\gamma$ as a function of $v_{2,{\rm single}}$ of single POIs selecting events on $\hat{q}_{2,{\rm pair}}^{2}$ of pairs of POIs reflects most closely the true CME in avfd. However, since the same POIs are used for $\hat{q}_{2,{\rm pair}}^{2}$, $v_{2,{\rm single}}$ and $\Delta\gamma$, self-correlations are present in the measurement whose effects/biases are hard to discern. It is unclear how much background is remaining in the ESS intercept, and whether the remaining background is positive or negative. This is illustrated by the cartoon in Fig. 1 (right). In this work, we focus on the best combination of pair $\hat{q}_{2,{\rm pair}}^{2}$ and single $v_{2,{\rm single}}$ in ESS as found in [36], and examine what possible biases there may exist and the possible level of remaining background in the ESS intercept.
@@ III.1 AVFD Model
The avfd model [50, 51, 52] is an anomalous fluid dynamics developed to describe the evolution of chiral fermion currents in the quark-gluon plasma (QGP) created in relativistic heavy ion collisions in addition to the normal VISHNU simulations [53]. The average magnetic field is calculated from the spectator protons, and it is nominally along the direction perpendicular to the RP. In the model, the magnetic field is quantified by event-by-event simulations [54] taking into account the fluctuations of the magnetic field direction with respect to the RP. A modest time evolution is assumed for the decreasing magnetic field with a typical lifetime comparable to the initial time of hydrodynamic evolution $\sim 0.6\text{ fm}/c$. The initial condition for the axial charge density ($n_{5}$) is dynamically generated in avfd to be proportional to the entropy density ($s$), and the strength is set via the proportionality coefficient $n_{5}/s$. Three $n_{5}/s$ values are used in the simulations, $n_{5}/s$=0, 0.1, and 0.2. The QGP medium evolution is simulated by viscous hydrodynamics to describe the bulk background in heavy ion collisions, with transport parameters for the diffusion coefficient as well as the relaxation time. A final hadronic stage after the hydrodynamic freeze-out is included with hadronic re-scatterings and resonance decays.
@@ III.5 Toy Model I
As default only 2% of the measured $K_{S}$ abundance are generated in our toy model because this is the approximate fraction of $K_{S}$ whose decay daughters are both reconstructed as primordial pions in the STAR experiment. We vary this fraction to investigate its effect on our results.
@@ IV.3 Toy Model Verification
We have conducted a systematic study using four physics models (avfd, ampt, epos4, and hydjet++) describing heavy ion collisions. It is found, with no CME signal, that the ESS intercept ($\Delta\gamma_{\textsc{ess}}$) can be negative, zero, or positive depending on the models and multiplicities of the events, and the ESE intercept ($\Delta\gamma_{\textsc{ese}}$) is mostly consistent with zero, as expected, albeit inherently large statistical uncertainties.
It is further found, using toy models, that $\Delta\gamma_{\textsc{ess}}$ appears to depend on the details of the event content, for example, the relative abundances of resonances and the presence of a mass continuum. Given that the event-shape selection variable $\hat{q}_{2,{\rm pair}}^{2}$, the single particle $v_{2,{\rm single}}$, and the three-point correlator $\Delta\gamma$ in $\hat{q}_{2,{\rm pair}}^{2}$-classified events all depend on the particles of interest (POIs), which are made of not only primordial particles but also all decay products of electromagnetic- and strong-decay resonances and partial products of weak-decay resonances, we conjecture that the ESS intercept $\Delta\gamma_{\textsc{ess}}$ is a complicated function of all elements that make up the event. These elements include resonance abundances relative to primordial pions, resonance $v_{2}$’s and $p_{\rm T}$ spectra, as well as primordial particle $v_{2}$ and $p_{\rm T}$ spectra. Depending on the detailed physics of these elements, $\Delta\gamma_{\textsc{ess}}$ can be positive, negative, or zero. The issue is, unfortunately, that it is unknown, needless to say proven, what combinations of all those elements in the event would give the ideal $\Delta\gamma_{\textsc{ess}}=0$ in the absence of CME. This makes it impractical to search for CME using the ESS method.

### v5

# Investigating event-shape methods in the search for the chiral magnetic effect in relativistic heavy ion collisions

@@ Abstract
The Chiral Magnetic Effect (CME) is a phenomenon in which electric charge is separated by a strong magnetic field from local domains of chirality imbalance and parity violation in quantum chromodynamics. The CME-sensitive observable, the charge-dependent three-point azimuthal correlator $\Delta\gamma$, is contaminated by a major physics background proportional to the particle’s elliptic flow anisotropy $v_{2}$. Event-shape engineering (ESE) binning events in dynamical fluctuations of $v_{2}$ and event-shape selection (ESS) binning events in statistical fluctuations of $v_{2}$ are two methods to search for the CME by projecting $\Delta\gamma$ to the measured anisotropy $v_{2}=0$ intercept. We conduct a systematic study of these two methods using physics models as well as toy model simulations. It is observed that the ESE method fulfills the general premise of measuring the CME but is statistically hungry. It is found that the intercept from the ESS method depends on the details of the event content, such as the mixtures of background-contributing sources, because of statistical fluctuations of intertwining variables used in the method, and is thus not practically useful to measure the CME.
@@ I Introduction
The focus of our study is placed on the effects of the event classifying variables on the intercept under various circumstances of the simulated events. We note, however, that the $v_{2}$ measured in experiments contains not only the flow anisotropy but also the so-called nonflow correlations, dependent on analysis methods. Nonflow correlations refer to those unrelated to the collision geometry common to the entire event, such as resonance decays and (di-)jet correlations [37, 38, 39, 40]. While we touch upon the nonflow issue in this article, a full and thorough discussion of nonflow effects on the ESE and ESS methods is outside the scope of the present study.
@@
II Event-Shape Techniques
@@
Since the ESE method selects events according to dynamical fluctuations of the $\left\langle v_{2}\right\rangle$ of the POIs, decoupled from the $q_{2}^{2}$ computation, it is guaranteed that the average resonance $\left\langle v_{\rm 2,{\rm res}}\right\rangle$, and generally the elliptic anisotropies of all elements in the event including all CME background sources, are proportional to $\left\langle v_{2}\right\rangle$. Note that the final-state $\left\langle v_{2}\right\rangle$ contains contributions not only from primordial particles, but also from all decay products of resonances and clusters. Since the averages $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ and $\left\langle v_{2}\right\rangle$ in those $q_{2}$-selected event classes are all connected to the initial geometry, they are all proportional to each other. This is illustrated in the cartoon in Fig. 1 (left), where the three ellipses illustrate event selections by $q_{2}$ and the hollow circles are the average $\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ in each $q_{2}$-selected event class. In the ESE analyses for CME searches [29, 30], one analyzes $\Delta\gamma$ as a function of $\left\langle v_{2}\right\rangle$ in each event class and projects the $\Delta\gamma$ measurement to zero $\left\langle v_{2}\right\rangle$. The ESE method guarantees that the elliptic flows of the background sources also vanish at zero $\left\langle v_{2}\right\rangle$. In other words, a linear projection to $\left\langle v_{2}\right\rangle=0$ would be equivalent to a projection to $\left\langle v_{\rm 2,{\rm res}}\right\rangle=0$, where the CME backgrounds are zero. Since all those events are from a narrow centrality bin where the CME signal does not vary much, the intercept of the linear projection is presumably the CME signal. Note that there may still be complications from nonflow effects [47] depending on how $\Delta\gamma$ and $\left\langle v_{2}\right\rangle$ are measured, which is outside the scope of the present work.
@@
Since the CME background comes from the elliptic anisotropies imprinted in the POIs, it is interesting to select events with an event-shape ellipticity $v_{2}^{\rm obs}\equiv\left\langle\cos 2(\phi-\psi)\right\rangle$ calculated by POIs themselves, where the event-plane (EP) azimuthal angle $\psi$ is determined in another kinematic region away from the POIs, and examine the $\Delta\gamma$ observable as a function of $v_{2}^{\rm obs}$. This method selects on the event-by-event elliptic-shape quantity (particle emissin pattern) $v_{2}^{\rm obs}$ which can fluctuate even to negative values because of large statistical fluctuations, the intercept at $v_{2}^{\rm obs}=0$, more sensitive to the CME, can be well determined from data. This event-shape (ES) analysis was performed by STAR [24] and the intercept of the linear fit is consistent with zero with the then-available statistics. However, the backgrounds are not determined by $v_{2}^{\rm obs}$ but by $\left\langle v_{\rm 2,{\rm res}}\right\rangle$. Statistically fluctuated $v_{2}^{\rm obs}=0$ does not necessarily guarantee $\left\langle v_{\rm 2,{\rm res}}\right\rangle=0$. This is illustrated by the cartoon in Fig. 1 (center). Although $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ is proportional to $\left\langle v_{2}\right\rangle$ with an unbiased event selection, the strongly (maximally) biased selection of events by $v_{2}^{\rm obs}$ will obscure the relationship between the averages $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ and the $v_{2}^{\rm obs}$ in each bin of $v_{2}^{\rm obs}$ of those events, as indicated by the hollow box in Fig. 1 (center). Because of the overall positive means $\left\langle v_{2}^{\rm obs}\right\rangle=\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ over the entire event sample, the linear projection will give a positive intercept of $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ at $v_{2}^{\rm obs}=0$. In other words, with finite positive average $\left\langle v_{2}\right\rangle$ in heavy ion collisions, $\left\langle v_{\rm 2,{\rm res}}\right\rangle>0$ at $v_{2}^{\rm obs}=0$ with events binned in $v_{2}^{\rm obs}$. This has been verified by model study [48]. Thus, positive residual background remains in the linear fit intercept in the analysis of $\Delta\gamma$ as a function of $v_{2}^{\rm obs}$ [24], although the background is significantly reduced compared to that in the inclusive $\Delta\gamma$ measurement. Ideally, in the ES analysis, one would want to select events with the $v_{2}^{\rm obs}$ values of all resonances (background sources) to be zero, however, such selection is experimentally insurmountable.
@@
It has been recently proposed by Xu et al. [36] to select events according to the $\hat{q}_{2,\textsc{poi}}^{2}$ of the POIs, and study the $\Delta\gamma$ as a function of $v_{2}$ of the POIs. This is similar in spirit to the ES method above where the events are selected by $v_{2}^{\rm obs}$ and the $\Delta\gamma$ is examined as a function of the same $v_{2}^{\rm obs}$ variable. Here, the selection variable $\hat{q}_{2,\textsc{poi}}^{2}$ and the $v_{2}$ variable are different, but related as they are computed by the same POIs. The $\hat{q}_{2}^{2}$ variable is an event-by-event quantity connected to two-particle cumulant anisotropy (cf Eq. 5), whereas $v_{2}$ can be measured with respect to $\psi$ from another kinematic region like in the ES method. (Of course $v_{2}$ can also be computed by the two-particle cumulant method, in which case it is more directly connected to $\hat{q}_{2}^{2}$.) The variations in the event-by-event $\hat{q}_{2,\textsc{poi}}^{2}$ quantity, within a given narrow centrality bin, are mainly from statistical fluctuations, just like the $\hat{q}_{2}^{2}$ in the ESE method. However, unlike ESE, the $v_{2}$ variable in those $\hat{q}_{2,\textsc{poi}}^{2}$-selected events by ESS are related to $\hat{q}_{2,\textsc{poi}}^{2}$ itself and are calculated from the same POIs, so the variations in $v_{2}$ is also primarily of statistical nature. (Note, we simply use $v_{2}$ here, instead of the $\left\langle v_{2}\right\rangle$ for ESE, to indicate that the $v_{2}$ in ESS is of statistical nature although it is an “average” $v_{2}$ in events selected by $\hat{q}_{2,\textsc{poi}}^{2}$.) This technique selects on the statistical fluctuations in event shape, not on the dynamical fluctuations in $\left\langle v_{2}\right\rangle$ like in the ESE method, and it is referred to as the Event-Shape-Selection (ESS) method [36]. The most important difference in the ESS method from the ESE method is that the $\hat{q}_{2}^{2}$ is computed from the POIs in the former but in the later the momentum space in which to compute $\hat{q}_{2}^{2}$ is displaced from that of the POIs’. In both approaches, one groups events (within a given narrow centrality bin) according to $\hat{q}_{2}^{2}$ and studies the $\Delta\gamma$ variable as a function of the $v_{2}$ of the POIs. It is worthy to emphasize again that, while the event selection variables in both methods are dominated by statistical fluctuations event-by-event, these statistical fluctuations are canceled in $v_{2}$ in the ESE method because the calculations of $\hat{q}_{2}^{2}$ and $v_{2}$ use different sets of particles, but do not cancel (rather strongly correlated) in the $v_{2}$ in the ESS method because the calculations of $\hat{q}_{2,{\rm pair}}^{2}$ and $v_{2}$ use the same POIs.
@@
One can study the $\Delta\gamma$ as a function of $v_{2}$ of single POI particles (or $v_{2,{\rm pair}}$ of pairs of POI particles) selecting events on $\hat{q}_{2,\textsc{poi}}^{2}$ of single POIs (or $\hat{q}_{2,{\rm pair}}^{2}$ of pairs of POIs). It is found by the Anomalous-Viscous Fluid Dynamics (avfd) model study that, out of the four combinations, it is the best to use the combination of pair $\hat{q}_{2,{\rm pair}}^{2}$ and single $v_{2,{\rm single}}$ [36], because the linear intercept of $\Delta\gamma$ as a function of $v_{2,{\rm single}}$ of single POIs selecting events on $\hat{q}_{2,{\rm pair}}^{2}$ of pairs of POIs reflects most closely the true CME in avfd. However, since the same POIs are used for $\hat{q}_{2,{\rm pair}}^{2}$, $v_{2,{\rm single}}$ and $\Delta\gamma$, self-correlations are present in the measurement the effects/biases of which are hard to discern. It is unclear how much background is remaining in the ESS intercept, and whether the remaining background is positive or negative. This is illustrated by the cartoon in Fig. 1 (right). In this work, we focus on the best combination of pair $\hat{q}_{2,{\rm pair}}^{2}$ and single $v_{2,{\rm single}}$ in ESS as found in Ref. [36], and examine what possible biases may exist and the possible level of remaining background in the ESS intercept.
@@ III.1 AVFD Model
The avfd model [50, 51, 52] is an anomalous fluid dynamics developed to describe the evolution of chiral fermion currents in the quark-gluon plasma (QGP) created in relativistic heavy ion collisions in addition to the normal simulations by VISHNU (a package combining the hydrodynamic evolution of the QGP with a microscopic hadronic cascade) [53]. The average magnetic field is calculated from the spectator protons, and it is nominally along the direction perpendicular to the RP. In the model, the magnetic field is quantified by event-by-event simulations [54] taking into account the fluctuations of the magnetic field direction with respect to the RP. A modest time evolution is assumed for the decreasing magnetic field with a typical lifetime comparable to the initial time of hydrodynamic evolution $\sim 0.6\text{ fm}/c$. The initial condition for the axial charge density ($n_{5}$) is dynamically generated in avfd to be proportional to the entropy density ($s$), and the strength is set via the proportionality coefficient $n_{5}/s$. Three $n_{5}/s$ values are used in the simulations, $n_{5}/s$=0, 0.1, and 0.2. The QGP medium evolution is simulated by viscous hydrodynamics to describe the bulk background in heavy ion collisions, with transport parameters for the diffusion coefficient as well as the relaxation time. A final hadronic stage after the hydrodynamic freeze-out is included with hadronic re-scatterings and resonance decays.
@@ III.5 Toy Model I
As default only 2% of the measured $K_{S}$ abundance are generated in our toy model because this is the approximate fraction of $K_{S}$ the decay daughters of which are both reconstructed as primordial pions in the STAR experiment. We vary this fraction to investigate its effect on our results.
@@ IV.3 Toy Model Verification
We have conducted a systematic study using four physics models (avfd, ampt, epos4, and hydjet++) describing heavy ion collisions. It is found, with no CME signal, that the ESS intercept ($\Delta\gamma_{\textsc{ess}}$) can be negative, zero, or positive depending on the models and multiplicities of the events, and the ESE intercept ($\Delta\gamma_{\textsc{ese}}$) is mostly consistent with zero, as expected, albeit with inherently large statistical uncertainties.
It is further found, using toy models, that $\Delta\gamma_{\textsc{ess}}$ appears to depend on the details of the event content, for example, the relative abundances of resonances and the presence of a mass continuum. Given that the event-shape selection variable $\hat{q}_{2,{\rm pair}}^{2}$, the single particle $v_{2,{\rm single}}$, and the three-point correlator $\Delta\gamma$ in $\hat{q}_{2,{\rm pair}}^{2}$-classified events all depend on the particles of interest (POIs), which are made of not only primordial particles but also all decay products of electromagnetic- and strong-decay resonances and partial products of weak-decay resonances, we conjecture that the ESS intercept $\Delta\gamma_{\textsc{ess}}$ is a complicated function of all elements that make up the event. These elements include resonance abundances relative to primordial pions, resonance $v_{2}$’s and $p_{\rm T}$ spectra, as well as primordial particle $v_{2}$ and $p_{\rm T}$ spectra. Depending on the detailed physics of these elements, $\Delta\gamma_{\textsc{ess}}$ can be positive, negative, or zero. The issue is, unfortunately, that it is unknown, and needless to say unproven, what combinations of all those elements in the event would give the ideal $\Delta\gamma_{\textsc{ess}}=0$ in the absence of CME. This makes it impractical to search for CME using the ESS method.